OpenStudy (anonymous):

Plzz help! - Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3.

OpenStudy (anonymous):

A. y = -3x - 13 B. y = -3x + 11 C. y = -3x + 13 D. y = -3x + 1

OpenStudy (anonymous):

Okay so this is the formula. y=mx+b. b is the y intercept and m is the slope. http://prntscr.com/6lr96n look at this link to find the y-intercept.

OpenStudy (anonymous):

Which one do you think it is?

OpenStudy (anonymous):

wait one sec

OpenStudy (anonymous):

could u help me out plz @GretaKnows

OpenStudy (anonymous):

Okay so plug in the numbers. m=-3. and b=1. y=-3x+1

OpenStudy (anonymous):

the y-intercept is where the line cross the y-axis.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so would the answer be... C?

OpenStudy (anonymous):

@GretaKnows

OpenStudy (anonymous):

wait sorry D?

OpenStudy (anonymous):

Yes.